The magnitude of the force of gravity on you, when you’re at rest on the surface of Earth, is your weight. How does your weight on Earth compare with your weight on Mars (which is the magnitude of the force of gravity on you, when you’re at rest on the surface of Mars)?
Answer #1

The gravitational force on you when you are at rest on earth is equal to weight = mg (the value of gravitational acceleration g on earth is approximately 9.8 m/s^2) so weight= (m)*( 9.8)

While being on mars as the weight =mg' (the value of gravitational acceleration g' on mars is approximately 3.7 m/s^2) so weight =(m)*(3.7)

Now by the values of both weights we can clearly observe that the weight on mars is less that weight on earth.

Dear student. feel Free to ask doubts in comment(give me a try before downvoting). Trust me ! I'll surely reply within few seconds and clear your doubt.


Hope you give thumbs up.
Thanks!